INAF queue observing, 2016 May 04-05

Observer: R. Carini, M. Di Criscienzo
Telescope Operator: D. Gonzalo-Huerta
Support Astronomer: B. Rothberg on call
Instrument: LBC, MODS, LUCI

Summary

At the sunset the sky is clear. wind velocity 10m/s During the flat fields the wind is increasing At 5.10 the TO closes the dome, the wind rear is 22 m/s At 8:00 the wind decreased, the TO opens the dome

Night duration (12\xB0 el twilight) 20:30-4:30 ( 03:30-11.30 UT) = 8 h

Observing time: * 5 h 10 m* (64%)

Weather time loss: * 2h 50m* (36%)

technical time loss * h* (00%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
FLAT LBC   I,z in the red channel    
Giallongo Virgo field 2 dual grating, rsloan 15 minutes completed the mods obs 3 and lbc ob 60s_90deg, seeing out of spec
mods bias        
standard S875-C_H2 H2,Ks        
Beltran A5 G28.87+0.07 H2, Ks 2h 15m completed

Details

02:59 acquisition standard star feige67
  • acqMODS feige67.acq mods1r.20160505.0002-0004
  • execMODS feige67_ls.obs mods1r.20160505.0005,mods1b.20160505.0002
  • seeing very variable between 1.2''-2''

03:15 acquisition science target, Giallongo's program, Virgo field2
  • acqMODS virgo2.acq mods1r.0006-0009
  • dofpia
  • 03:42 execMODS mods1r.0010,mods1b.0003
  • run virgo_rsloan_60s_90deg.ob
  • seeing is still variable, 1.3''-2.0''
  • 03:50 some clouds are coming

04:00 the seeing is too bad, Avg FWHM 1.8'', we switch to LUCI2 for a backup program

04:41 begin the acquisition of the target Set2_1ControlJ110024.57+320218.7 of the the set2 of the Beltran's program
  • luci2.20160505.0001-0026 (wrong number of Nexp during the manually offset)
  • mean seeing 1.8
  • some cirrus are present
  • the wind is increasing

05:10 the to closes the dome, the wind is too strong. wind rear 22m/s

* bias mods mods1r.0011-0020,mods1b.0003-0013

* ./doBias

08:00 the TO opens the dome. The wind has decreased, wind rear is 16m/s, the sky is clear

* we begin with the Beltran's program, acquire the telluric star
  • GO S875-C_H2.img.xml luci2.20160505.0027-0031
  • seeing variable 1.2''-1.6''
  • are present again the obscuration at the bottom of the field
  • GO S875-C_Ks.img.xml luci2.20160505.0032-0036
  • seeing 1.0''

08:47 slew to scientific target A5 luci2.20160505.0037-0066
  • GO A5_H2.img.xml
  • seeing 0.9''-1.0''
  • 09:00 problem with the guider due to the wind gusts. bad seeing =1.3''
  • 09:05 the seeing is variable between 1.0 and 1.5'' *09:46 the seeing is more stable about 1.0'' * 10:15 seeing 0.9''

10:25 due to initial variable seeing we take other images for this target in H2 luci2.20160505.0067-0076
  • seeing 0.9''
  • 10:44 seeing 0.8''
  • 10:52 seeing 0.9''-1.0''

10:57 GO A5_Ks.img.xml luci2.20160505.0077-0096
  • seeing 0.9''-1.0''

11:36 we take again data for the standard in H2 target with better seeing
  • seeing 1.0''
  • acquisition during the sunset

-- %USERSIG{RobertaCarini - 2016-05-05}%

Comments

 
Topic revision: r6 - 06 May 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback